How many aluminum atoms are in 4.22 g of aluminum?


1
Expert's answer
2022-04-11T00:33:03-0400

N = n * NA

n(Al) = m(Al)/M(Al)

n(Al) = 4.22 g / 27g/moll = 0.156 moll

N = 0.156 * 6.02*1023 = 9.39*1023 atoms